# Central limit theorems for interacting innovation processes, related statistical tools and general results

### v3

@@ Abstract.
Understanding the mechanisms by which novelties emerge, propagate, and catalyze further innovations is pivotal across various disciplines, including biology, linguistics, and social sciences Armano and Javarone (2017); Arthur (2009); Fink et al. (2017); Puglisi et al. (2008); Rogers et al. (2019); Rzhetsky et al. (2015)). In probabilistic terms, a novelty or innovation is defined as the first occurrence of an event of interest.
A widely adopted mathematical model for such innovation process is an urn model with infinitely many colors Bandyopadhyay and Thacker (2022); Sariev et al. (2023), also known as species sampling sequence Favaro et al. (2009); Hansen and Pitman (2000); Lijoi et al. (2007); Lijoi and Prünster (2010); Pitman (1996). For each $t\in{\mathbb{N}}\setminus\{0\}$, let $C_{t}$ denote the observed color at extraction $t$. The probability that the $(t+1)$-th draw yields a “new” color (i.e., one not previously observed) is denoted by $Z^{*}_{t}$ and it is a function of the observations $C_{1},\dots,C_{t}$. Conversely, the probability that the next color is an “old” (i.e., already observed) color $c$ is given by $P_{c,t}=\sum_{n=1}^{t}Q_{n,t}\mathbb{I}_{\{C_{n}=c\}}$, where $Q_{n,t}$ depends on the observations $C_{1},\dots,C_{t}$. The quantities $Z^{*}_{t}$ and $Q_{n,t}$ specify the model: precisely, $Z^{*}_{t}$ represents the birth probability, indicating the likelihood of encountering a new color (that is a novelty) at time-step $t+1$, while $Q_{n,t}$ signifies the weight assigned to the $n$-th extraction at time-step $t$, reflecting a reinforcement mechanism where the probability of drawing a particular color increases with its past occurrences.
In this framework, the set of possible colors is unbounded, allowing for the continuous introduction of new colors. The urn with infinitely many colors symbolizes the space of all potential events, whereas the sequence of extracted balls represents the realized history. The Poisson-Dirichlet (PD) model James et al. (2008); Pitman (2006); Pitman and Yor (1997) serves as a canonical example of an innovation process. In this model we have
@@
where $0\leq\gamma<1$, $\theta>-\gamma$, $D_{t}$ denotes the number of distinct colors observed up to time-step $t$ and $K_{c,t}$ is the count of color $c$ up to time-step $t$. This model embodies the “preferential attachment” or “popularity” principle, wherein the probability of observing a particular color is simply proportional to its frequency. The PD process effectively captures empirical laws such as Heaps’ law (Heaps, 1978), which describes the sublinear growth of distinct elements with respect to the total number of observations, and Zipf’s law (Zipf, 1949), characterizing the frequency-rank distribution of elements. Specifically, in the PD model the number of distinct elements $D_{t}$ scales as $D_{t}\propto t^{\gamma}$ for $0<\gamma<1$, and $D_{t}\propto\ln(t)$ when $\gamma=0$.
@@
An intuitive representation of the PD process is the urn with triggering model Aletti and Crimaldi (2021); Tria et al. (2014). An urn initially contains $N_{0}>0$ distinct balls of different colors. Then, at each time-step $t+1$, a ball is drawn at random with replacement and the urn is updated as follows:
@@
  • if the drawn color is new, i.e. not previously observed, then we add $(\widehat{\rho}-1)$ balls of the same color plus $(\nu+1)$ balls of entirely new (not already present in the urn), distinct colors;
  • if the drawn color is old, i.e. it has been already observed, we add $\rho$ balls of the same color.
@@
When the balance condition $\widehat{\rho}+\nu=\rho$ is satisfied, ensuring a constant number of balls added at each time-step, the model reproduces the probabilities specified in equation (0.1) by taking $\rho>\nu\geq 0$ and setting $\theta=N_{0}/\rho$ and $\gamma=\nu/\rho$.
This formulation of the PD process highlights not only the reinforcement mechanism, but also the Kauffman’s principle of the adjacent possible (Kauffman, 2000), wherein the occurrence of a novelty expands the space of potential future novelties. Indeed, the model starts with an urn with a finite number of balls with distinct colors and, whenever a color is extracted for the first time, a set of balls with new colors is added to the urn. This represents Kauffman’s idea that, when a novelty occurs, it triggers further potential novelties.
@@
To explore how different innovation processes influence each other, in Aletti et al. (2023) we have introduced and analyzed a model for a finite network of urns with triggering (see Section 1). Each node in the network possesses its own urn, and the probability of observing a new or old item at a certain node depends on both the node’s own history and the histories of the other nodes. The interactions are governed by two matrices: $\Gamma$ and $W$. Matrix $\Gamma$ modulates the influence of novelties produced in the system on the probability of each node to generate new novelties, extending the Kauffman’s principle of the adjacent possible to the system-level. Matrix $W$ regulates the reinforcement mechanism across nodes, letting the probability of drawing already observed items to be dependent on their frequencies in all the nodes.
@@
When the interaction matrices $\Gamma$ and $W$ are irreducible, the model exhibits synchronization phenomena. Specifically, all nodes produce innovations at the same asymptotic rate, and the distribution of the observed items becomes uniform across the network. These results are formalized in the following Theorems 1.1 and 1.2.
@@
  • derive the second-order asymptotic behaviour of the model, providing Central Limit Theorems (CLTs) for the quantities of interest;
  • develop statistical tools to make inference on the interaction matrices $\Gamma$ and $W$, as well as on the limiting probabilities of observing an extracted item $c$;
  • apply the theoretical findings to real-world datasets, assessing the interaction intensities in empirical contexts.
It is worthwhile to note that we prove general theoretical results, that can be applied also in other frameworks. Indeed, they regard a pair of multidimensional stochastic processes following a Stochastic Approximation (SA) dynamics (e.g. Laruelle and Pagès (2013); Mokkadem and Pelletier (2006)). However, CLTs from SA literature do not apply to our situation. Indeed, the asymptotic covariance matrices in those results are deterministic, while in our CLTs they are random. This is why we do not use the simple convergence in distribution, but we employ the notion of stable convergence, which is, among other things, essential for obtaining the statistical tools.
@@ 0.1. Related literature
While extensive literature exists on interacting urn models with a finite set of colors (see, for instance, Aletti et al. (2017, 2024) and the references therein), studies focusing on urns with infinitely many colors are relatively scarce. Notable exceptions include (Fortini et al., 2018; Iacopini et al., 2020), which explore Dirichlet processes with random reinforcement and urns with triggering, respectively. However, these models differ in key aspects, such as the definition of novelty and the nature of interactions, distinguishing our approach in its system-level application of the adjacent possible principle and the dual reinforcement mechanism. In the model provided in Fortini et al. (2018) (see Example 3.8 in that paper), there is a finite collection of Dirichlet processes with random reinforcement. More precisely, in that model we have a random weight $W_{t,h}$ associated to the extraction at time-step $t$ from the urn $h$ so that, the probability of extracting from urn $h$ an old color (here, the term “old” is not at the system-level, but refers to urn $h$ only) is proportional to the weight associated to that color. The interaction across the urns is introduced by means of the weights, which could be stochastically dependent: each $W_{t,h}$ may be the same for each urn $h$, or a function of the observed outcomes of the other urns, or a function of some common (observable or latent) variables. The present work and Iacopini et al. (2020) share the fact that the considered models are both a collection of urns with triggering with an interacting dynamics that brings the Kauffman’s principle of the adjacent possible from the single agent to the network of agents. Adopting the terminology of Iacopini et al. (2020), we can say that both interacting mechanisms are based on the construction and the updating of a “social” urn for each network node from which the extractions take place, but the construction and the updating rules of the social urns are deeply different in the two models. In particular, differently from Iacopini et al. (2020), we introduce the notion of “new” and “old” at the system level. In Iacopini et al. (2020) the authors focus on the novelties in each sequence (novelty in $h$ = first appearance in $h$ of a new item), that they call “discoveries”; while we also study the sequence of the novelties for the whole system produced by each agent. Furthermore, in Iacopini et al. (2020) the extraction of an “old” item in a certain network node does not affect the other nodes, in our model we also have an interacting reinforcement mechanism for the “old” items: the probability of the extraction of an old item depends on the number of times it has been observed in all the nodes. This allows us to get a specific result on the distribution of the observations in the system among the different items observed.
@@ 0.2. Structure of the work
The remainder of the paper is organized as follows. In Section 1 we describe the model and we recall the first-order convergence results. In Section 2 we present the main results of this paper, which consist in the CLTs describing the second-order asymptotic behaviour of the considered system of interacting innovation processes. Then, in Section 3 we illustrate statistical tools derived from the CLTs. To facilitate its application across a wide range of contexts, the inferential methodology is presented within a general framework in the separate appendix, which also includes the general theoretical results and their proofs. Sections, theorems and remarks reported in the appendix are referenced with capital letters. Finally, in Section 4 we apply the proposed methods to two real data sets, evaluating the interaction intensities.
@@ 1. Model and first-order convergence results
Suppose to have $N$ urns, labeled from $1$ to $N$. At time-step 0, the colors inside each urn are different from those in the other urns. Let $N_{0,h}>0$ be the number of distinct balls with distinct colors inside the urn $h$. Then, at each time-step $t\geq 1$, one ball is randomly drawn with replacement from each urn and, for any $h=1,\dots N$, urn $h$ is so updated according to the colors extracted from urn $h$ itself and from all the other urns $j\neq h$:
@@
The terms “new” and “old” refer to the entire system, that is a “new” color is a color that has never been extracted from an urn of the system. On the contrary, an “old” color is a color that has already been extracted from at least one urn of the system. We assume that the “new” colors added to a certain urn are always different from those added to the other urns (at the same time-step or in the past). By means of this fact, together with the assumption that initially the colors in the urns are different from each other, we cannot have the same new color extracted simultaneously from different urns. In other words, we cannot have the same novelty observed simultaneously in different processes of the system. Therefore, for each observed new color (novelty) $c$, there exists a unique urn (process), say $j^{*}(c)$, in the system that produced it. However, in a time-step following its first extraction, color $c$ could be also extracted from another urn $h\neq j^{*}(c)$, as a consequence of the interaction among the urns (processes). Indeed, the “contamination” of the color-set of the urn $h$ with the colors present in the other urns is possible by means of the interaction terms $\widehat{\rho}_{j,h}$ and/or $\rho_{j,h}$ in the above model dynamics. Furthermore, as in the standard Poisson-Dirichlet model, we assume the balance condition
@@
where $K_{t}(j,c)$ denotes the number of times the color $c$ has been extracted from urn $j$ until time-step $t$ and $j^{*}(c)$ denotes the urn from which the color $c$ has been extracted for the first time. Note that the probability that urn (process) $h$ will produce at time-step $t+1$ a novelty for the entire system has an increasing dependence on the number $D^{*}_{t,j}$ of novelties produced by the urn (process) $j$ until time-step $t$ and the parameter $\gamma_{j,h}$ regulates this dependence. In other words, Kauffman’s principle of the adjacent possible is at the “system level”: that is, for each pair $(j,h)$ of urns in the system, the parameter $\gamma_{j,h}$ quantifies how much the production of a novelty by urn $j$ induces potential novelties in urn $h$. Moreover, on the other hand, the probability that from urn $h$ we will extract at time-step $t+1$ an old color $c$ has an increasing dependence on the number $K_{t}(j,c)$ of times the color $c$ has been drawn from urn $j$ until time-step $t$ and the parameter $w_{j,h}$ quantifies how much the number $K_{t}(j,c)$ leads toward a future extraction of a ball of color $c$ from urn $h$. In Aletti et al. (2023) we proved the first-order asymptotic properties of the model:
Theorem 1.1.
@@
Suppose that the matrix $\Gamma=(\gamma_{j,h})_{j,h=1,\dots,N}$ is irreducible and denote by $\gamma^{*}\in(0,1)$ the Perron-Frobenius eigenvalue of $\Gamma$. Then, for each $h=1,\dots,N$, we have
(1.5) $t^{-\gamma^{*}}D^{*}_{t,h}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}D^{**}_{\infty,h}\,$,
@@
where $D^{**}_{\infty,h}$ is a finite strictly positive random variable. Moreover, we have
$\frac{D^{*}_{t,h}}{D^{*}_{t,j}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\frac{u_{h}}{u_{j}}\quad\mbox{and}\quad\frac{D^{*}_{t,h}}{\sum_{j=1}^{N}D^{*}_{t,j}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\frac{u_{h}}{\sum_{j=1}^{N}u_{j}}\,$,
where ${\mathbf{u}}=(u_{h})_{h}$ is the (up to a multiplicative constant different from zero) left eigenvector of $\Gamma$ associated to $\gamma^{*}$.
Theorem 1.2.
@@
Suppose that the matrix $W=(w_{j,h})_{j,h=1,\dots,N}$ is irreducible. Then, for each $h=1,\dots,N$ and each observed color $c$ in the system, we have
(1.6) $P_{t}(h,c)\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\widetilde{P}_{\infty}(c)\,$,
@@
(1.7) $\frac{1}{t}K_{t}(h,c)\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\widetilde{P}_{\infty}(c)\quad\mbox{and}\quad\frac{K_{t}(h,c)}{\sum_{j=1}^{N}K_{t}(j,c)}\longrightarrow\frac{1}{N}\,$,
@@
where $\widetilde{P}_{\infty}(c)$ is a random variable that takes values in $(0,1)$ and does not depend on $h$  [^1] Note that in Th 3.2 of Aletti et al. (2023) we state that this random variable takes values in $(0,1]$, but, applying (Aletti et al., 2023, Theorem S1.3) also replacing $P_{t}(h,c)$ by $1-P_{t}(h,c)$, we immediately obtain $(1-\widetilde{P}_{\infty}(c))>0$ a.s., that is $\widetilde{P}_{\infty}(c)<1$ a.s...
The first result states that, if $\Gamma$ is irreducible, that is the graph with the innovation processes as nodes and with $\Gamma$ as the adjacency matrix is strongly connected, then $D^{*}_{t,h}\propto t^{\gamma^{*}}$ a.s. for all $h=1,\dots,N$, that is all the $D^{*}_{t,h}$ grow with the same Heaps’ exponent $\gamma^{*}\in(0,1)$. This means that, at the steady state, all the processes of the network produce innovations for the system at the same rate. In addition, the ratio $D^{*}_{t,h}/D^{*}_{t,j}$ provides a strongly consistent estimator of the ratio $u_{h}/u_{j}$ of the relative eigenvector centrality scores (with respect to $\Gamma^{\top}$) of the two nodes $h$ and $j$ and the share of innovations produced by process $h$ for all the system converges almost surely to the absolute eigenvector centrality score of $h$. The second result affirms that if $W$ is irreducible, that is the graph with the innovation processes as nodes and $W$ as the adjacency matrix is strongly connected, then, for each observed item $c$, the number of times item $c$ has been observed in process $h$ grows linearly. Moreover, the number of times item $c$ has been observed is asymptotically uniformly distributed among the processes of the system and, at the steady state, the probability $\widetilde{P}_{\infty}(c)$ that an item $c$ is observed in a process is the same for each process.

### v4

@@ Abstract.
1. Introduction
Understanding how novelties emerge, propagate, and catalyze further innovations is crucial in several disciplines, including biology, linguistics, and social sciences Armano and Javarone, (2017); Arthur, (2009); Fink et al., (2017); Puglisi et al., (2008); Rogers et al., (2019); Rzhetsky et al., (2015). In probabilistic terms, a novelty (or innovation) is the first occurrence of an event of interest.
Widely adopted mathematical models for such innovation processes include urn models with infinitely many colors Bandyopadhyay and Thacker, (2022); Mailler and Villemonais, (2020); Sariev et al., (2023); Sariev and Savov, (2024) and species sampling sequences Favaro et al., (2009); Hansen and Pitman, (2000); Lijoi et al., (2007); Lijoi and Prünster, (2010); Pitman, (1996). In the urn metaphor, for each $t\in{\mathbb{N}}\setminus\{0\}$ we denote by $C_{t}$ the observed color at extraction $t$. The probability that the $(t+1)$-th draw yields a “new” color (i.e., one not previously observed) is $Z^{*}_{t}$, a function of $C_{1},\dots,C_{t}$, while the probability of drawing an “old” (i.e., already observed) color $c$ is $P_{c,t}=\sum_{n=1}^{t}Q_{n,t}\,\mathbb{I}_{\{C_{n}=c\}}$, where $Q_{n,t}$ depends on $C_{1},\dots,C_{t}$. The quantities $Z^{*}_{t}$ and $Q_{n,t}$ specify the model. In this framework, the set of possible colors is unbounded, so new colors may continuously enter the system: the urn represents the space of possibilities, and the extracted sequence the realized history.
A canonical example is the Poisson-Dirichlet (PD) model James et al., (2008); Pitman and Yor, (1997). In this model,
@@ 1. Introduction
where $0\leq\gamma<1$, $\theta>-\gamma$, $D_{t}$ is the number of distinct colors observed up to time-step $t$, and $K_{c,t}$ is the count of color $c$ up to time-step $t$. This model embodies the “preferential attachment” (or “popularity”) principle, since the probability of observing a color is proportional to its frequency. The PD process effectively captures empirical laws such as Heaps’ law (Heaps,, 1978), describing the sublinear growth of distinct elements with the number of observations, and Zipf’s law (Zipf,, 1949), characterizing the sublinear decrease of ranked frequencies. In particular, in the PD model $D_{t}\propto t^{\gamma}$ for $0<\gamma<1$, and $D_{t}\propto\ln(t)$ when $\gamma=0$ Janson, (2005); Tria et al., (2018).
@@
An intuitive representation of the PD process is the urn with triggering model Aletti and Crimaldi, (2021); Tria et al., (2014) (or the equivalent formulation in Zabell, (2005)). Initially, the urn contains $N_{0}>0$ distinct balls of different colors. At each time-step $t+1$, a ball is drawn at random with replacement and:
@@
  • if the drawn color is new, we add $(\widehat{\rho}-1)$ balls of the same color and $(\nu+1)$ balls of entirely new (distinct) colors;
  • if the drawn color is old, we add $\rho$ balls of the same color.
@@
When the balance condition $\widehat{\rho}+\nu=\rho$ holds (so that the number of added balls is constant), the model reproduces (1.1) by taking $\rho>\nu\geq 0$ and setting $\theta=N_{0}/\rho$ and $\gamma=\nu/\rho$.
This formulation highlights both reinforcement and Kauffman’s principle of the adjacent possible (Kauffman,, 2000): when a novelty occurs, it expands the space of potential future novelties by introducing new colors.
@@
To study how distinct innovation processes influence each other, in Aletti et al., (2023) we introduced and analyzed a finite network of urns with triggering (see Section 2). Each node has its own urn, and the probability of observing a new or old item at a node depends on both the node’s history and the histories of the other nodes. Interactions are governed by two matrices: $\Gamma$ and $W$. Matrix $\Gamma$ modulates how novelties produced in the system affect each node’s probability of generating new novelties, extending the adjacent possible to the system level. Matrix $W$ regulates reinforcement across nodes, making the probability of extracting already observed items depend on their frequencies in all nodes.
@@
When $\Gamma$ and $W$ are irreducible, the model captures synchronization: all nodes produce innovations at the same asymptotic rate, and the distribution of observed items becomes uniform across the network. These results are formalized in Theorems 2.1 and 2.2.
@@
  • derive second-order asymptotic behaviour of the model by providing Central Limit Theorems (CLTs) for the quantities of interest;
  • develop statistical tools for inference on the interaction matrices $\Gamma$ and $W$, as well as on the limiting probabilities of observing an extracted item $c$;
  • apply the theoretical findings to real-world datasets, assessing interaction intensities in empirical contexts.
We also prove general theoretical results that apply beyond the present setting. They concern a pair of multidimensional stochastic processes following a suitable recursive dynamics. Since the asymptotic covariance matrices in our results are random, we do not rely on the standard convergence in distribution, but instead we employ the notion of stable convergence.
@@ 1.1. Related literature
Although interacting urn models with a finite set of colors are widely studied (see, e.g., Aletti et al., (2017, 2024) and references therein), works on urns with infinitely many colors are comparatively scarce. Notable exceptions are (Fortini et al.,, 2018; Iacopini et al.,, 2020), which investigate Dirichlet processes with random reinforcement and urns with triggering, respectively. These approaches differ from ours in the definition of novelty and in the interaction structure, in particular concerning the system-level adjacent possible and the dual reinforcement mechanism.
In Fortini et al., (2018) (see Example 3.8 therein), a finite collection of Dirichlet processes with random reinforcement is considered. A random weight $W_{t,h}$ is associated with the extraction at time-step $t$ from urn $h$, and the probability of extracting an old color (where “old” is defined at the level of urn $h$ only) is proportional to the weight of that color. Interactions across urns are introduced through stochastic dependence among the weights: for instance, the weights may be shared across urns, depend on outcomes of other urns, or depend on common observable or latent variables.
The present work and Iacopini et al., (2020) both study interacting urns with triggering, where interaction acts when a new color is extracted, thus extending Kauffman’s principle from a single agent to a network. In contrast with Iacopini et al., (2020), in our model the extraction of an “old” item is also interactive: the probability of extracting an old item depends on its total number of occurrences across all nodes. Moreover, Iacopini et al., (2020) focuses on node-level novelties (“discoveries”), i.e. first appearances in a given node, possibly already observed in other nodes, whereas we study the sequence of novelties for the whole network produced by each node.
@@ 1.2. Structure of the work
The remainder of the paper is organized as follows. In Section 2 we describe the model and recall the first-order convergence results. In Section 3 we present the main results, consisting of CLTs describing the second-order asymptotic behaviour of the interacting innovation processes. Then, in Section 4 we illustrate statistical tools derived from the CLTs. The inferential methodology is presented for a suitable bi-dimensional recursive process in a separate appendix, including theoretical results and proofs; sections, theorems, and remarks in the appendix are referenced with capital letters. Finally, in Section 5 we apply the proposed methods to two real datasets, evaluating interaction intensities.
@@ 2. Model and first-order convergence results
Suppose we have $N$ urns, labeled from $1$ to $N$. At time-step 0, the colors inside each urn are different from those in the other urns. Let $N_{0,h}>0$ be the number of distinct balls with distinct colors inside the urn $h$. Then, at each time-step $t\geq 1$, one ball is randomly drawn with replacement from each urn and, for any $h=1,\dots N$, urn $h$ is so updated according to the colors extracted from urn $h$ itself and from all the other urns $j\neq h$:
@@
The terms “new” and “old” refer to the entire system. We assume that the “new” colors added to a given urn are always different from those added to the other urns (at the same time-step or in the past). Together with the initial condition, this implies that the same novelty cannot be observed simultaneously in different urns. Hence, for each observed novelty $c$, there exists a unique urn, denoted $j^{*}(c)$, that produced it. After its first extraction, $c$ may subsequently appear in other urns due to interactions among urns. Furthermore, as in the standard Poisson-Dirichlet model, we assume the balance condition
@@
where $K_{t}(j,c)$ denotes the number of times the color $c$ has been extracted from urn $j$ until time-step $t$ and $j^{*}(c)$ denotes the urn from which the color $c$ has been extracted for the first time. Note that the probability that urn $h$ produces at time-step $t+1$ a novelty for the entire system depends increasingly on the number $D^{*}_{t,j}$ of novelties produced by urn $j$ up to time $t$, and $\gamma_{j,h}$ regulates this dependence. Thus, Kauffman’s principle of the adjacent possible holds at the system level: for each pair $(j,h)$, $\gamma_{j,h}$ quantifies how the production of a novelty by urn $j$ induces potential novelties in urn $h$. Similarly, the probability that urn $h$ extracts an old color $c$ at time-step $t+1$ depends increasingly on the counts $K_{t}(j,c)$, with $w_{j,h}$ quantifying the influence of urn $j$ on urn $h$ through reinforcement.
In Aletti et al., (2023) (see also (Aletti et al.,, 2025, Appendix B.1) for a general theory) we proved the first-order asymptotic properties of the model:
Theorem 2.1 ((Aletti et al.,, 2023, Theorem 3.1)).
@@
Suppose that the matrix $\Gamma=(\gamma_{j,h})_{j,h=1,\dots,N}$ is irreducible. Denote by $\gamma^{*}\in(0,1)$ the Perron-Frobenius eigenvalue of $\Gamma$ and by $\mathbf{u}$} its left eigenvector. Then, we have
$t^{-\gamma^{*}}{\mathbf{D}}^{*}_{t}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}D^{**}_{\infty}{\mathbf{u}}\,$,
@@
where ${\mathbf{D}}^{*}_{t}=(D^{*}_{t,1},\ldots,D^{*}_{t,N})^{\top}$ and $D^{**}_{\infty}$ is an integrable strictly positive random variable.
Theorem 2.2 ((Aletti et al.,, 2023, Theorem 3.2)).
@@
Suppose that the matrix $W=(w_{j,h})_{j,h=1,\dots,N}$ is irreducible. Then, for each observed color $c$ in the system, we have
${\mathbf{P}}_{t}(c)\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\widetilde{P}_{\infty}(c){\boldsymbol{1}}$
@@
$\frac{1}{t}{\mathbf{K}}_{t}(c)\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\widetilde{P}_{\infty}(c){\boldsymbol{1}}\quad\mbox{and so}\quad\frac{K_{t}(h,c)}{\sum_{j=1}^{N}K_{t}(j,c)}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\frac{1}{N}\,$,
@@
where ${\mathbf{P}}_{t}(c)=(P_{t}(1,c),\dots,P_{t}(N,c))^{\top}$, ${\mathbf{K}}_{t}(c)=(K_{t}(1,c),\dots,K_{t}(N,c))^{\top}$, $\boldsymbol{1}$}} denotes the vector with all the components equal to $1$ and $\widetilde{P}_{\infty}(c)$ is a random variable that takes values in $(0,1)$.  [^1] Note that in (Aletti et al.,, 2023, Theorem 3.2 ) we state that this random variable takes values in $(0,1]$, but, applying (Aletti et al.,, 2023, Theorem S1.3) also replacing $P_{t}(h,c)$ by $1-P_{t}(h,c)$, we immediately obtain $(1-\widetilde{P}_{\infty}(c))>0$ a.s., that is $\widetilde{P}_{\infty}(c)<1$ a.s..